OpenStudy (anonymous):

would (x+1)(x+1)=4x^2 ?

OpenStudy (anonymous):

2x+2

OpenStudy (anonymous):

so no

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

no it would equal \[x^2 +x +x +1 = x^2 +2x+1\]

OpenStudy (anonymous):

(x+1)(x+1) = x^+2x +1

OpenStudy (anonymous):

you add your two like terms, which equals 2x and then add your constants(which is 2)

OpenStudy (anonymous):

what r the 2 like terms?

OpenStudy (anonymous):

never mind i forgot to multiply but like terms r (x) bcuz they both end in ( x) so 4x is a like term with x but x^2 and 4x r not like terms

OpenStudy (anonymous):

oh, ok

myininaya (myininaya):

\[x^2+2x+1=4x^2\] \[0=4x^2-x^2-2x-1\] \[0=3x^2-2x-1\] \[0=3x^2-3x+x-1\] \[0=3x(x-1)+1(x-1)\] \[0=(x-1)(3x+1)=> x=1 \text{ or } x=\frac{-1}{3}\] so \[(x+1)(x+1)=4x^2 \text{ when } x=1 \text{ or } x=-\frac{1}{3}\]

OpenStudy (anonymous):

@myininaya, what does that have 2 do with my problem?

myininaya (myininaya):

its when those two expressions are equal

OpenStudy (anonymous):

what 2 expressions?

myininaya (myininaya):

4x^2 and (x+1)(x+1)

OpenStudy (anonymous):

@myininaya thts what i said but u wrote it out and did somesort of mathy fun stuff

myininaya (myininaya):

lol

OpenStudy (anonymous):

how r they equal?

myininaya (myininaya):

they are only equal when x=1 or x=-1/3 they are not equal for any other numbers

OpenStudy (anonymous):

why is that?

myininaya (myininaya):

4x^2 and (x+1)(x+1) do not intersect anywhere else

OpenStudy (anonymous):

they have the same variable

OpenStudy (anonymous):

@myininaya. what do u mean by intersceting any where else?

myininaya (myininaya):

their graphs only meet at x=1 or x=-1/3

myininaya (myininaya):

you can give a counterexample of why (x+1)(x+1)=4x^2 do not equal for all x if x=0, (x+1)(x+1)=(0+1)(0+1)=1(1)=1 4x^2=4(0)^2=4(0)=0 1 does not equal 0

OpenStudy (anonymous):

i'm confused at where u r heading

myininaya (myininaya):

you asked why they only equal for two values of x when we moved everything to one side we had a 2nd degree polynomial which means there are two solutions and both of those solutions were real (we got two different real solutions) so their graphs only interest in two different spots which were x=1 or x=-1/3 as showed above

myininaya (myininaya):

intersect*

OpenStudy (anonymous):

when u r talking about graphs and interseceting and stuff, what does that mean?

myininaya (myininaya):

you haven't heard of a graph?

OpenStudy (anonymous):

well, i have when it comes 2 comparing stuff and i did something similar 2 that in math class a week ago, so no not really

myininaya (myininaya):